# Exploration and Incentivized Participation in Randomized Trials

### v10

@@ 2 Our model: incentivized RCT (homogeneous agents)
We focus on homogeneous agents, i.e., agents with the same “type”, as stated in Section 1.2. We consider a randomized controlled trial (RCT) with some patients/users, respectively interpreted as a mechanism and agents. The mechanism interacts with the agents as a multi-armed bandit algorithm, but optimizes a very different statistical objective, and operates under economic constraints to incentivize participation. Accordingly, we split the model description into three subsections: the bandit model for the interaction protocol the statistical model for its performance objective, and the economic model for the agents’ incentives.
@@ 2.2 Statistical model: the performance objective
The mechanism should estimate the average score of each arm without any assumptions on the adversary, i.e., on how the outcomes are generated. Not relying on modeling assumptions is a standard desiderata in randomized controlled trials.
@@
We focus on the following two scenarios. First, $f_{\mathtt{adv}}(a)$ is the frequency of a given outcome $\omega^{*}\in\Omega$ in the main stage; then the scoring function is $f(\omega)={\bf 1}_{\left\{\,\omega=\omega^{*}\,\right\}}$. Second, if $f(\omega)$ is the outcome’s value (for the agent and/or the society), then $f_{\mathtt{adv}}(a)$ is the average value of arm $a$ [^16] In particular, the average treatment effect can be expressed as $f_{\mathtt{adv}}(\text{treatment})-f_{\mathtt{adv}}(\text{control})$. We keep $f$ abstract throughout to handle both scenarios in a uniform way.
@@ 2.3 Economic model: agents, beliefs and incentives
The mechanism must incentivize participation, i.e., satisfy Individual Rationality (IR). We require Bayesian IR in expectation over the agents’ beliefs over the adversaries. [^17] We invoke agents’ beliefs because satisfying the IR property for every adversary (a.k.a. ex-post IR) is clearly impossible, except for a trivial mechanism that always chooses the agent’s outside option. Crucially, we posit that agents’ beliefs are over stochastic adversaries, i.e., the agents believe that a given treatment yields the same outcome distribution in all agents. [^18] Agents have misspecified beliefs, in the sense that their support does not necessarily include the actual adversary. In particular, the actual adversary need not be stochastic. Misspecified beliefs are studied in a considerable recent literature in economic theory (starting from Berk,, 1966; Esponda and Pouzo,, 2016). However, this literature tends to focus on convergence and equilibria (e.g., Fudenberg et al.,, 2021; Esponda and Pouzo,, 2021; Esponda et al.,, 2021), whereas our paper takes the mechanism design perspective and is concerned with IPS-style statistical estimation. All agents share the same belief, a common simplification in economic theory. Such beliefs represent the current medical knowledge relevant to the clinical trial, or the learnings from user interactions collected by the online platform (which the platform may choose to share with its users).
@@ 2.4 Preliminaries
Per-round recommendation policy. A single round $t$ of our mechanism can be interpreted as a stand-alone “recommendation policy” which inputs the $\mathtt{hist}_{t}$, and outputs a recommended arm. Such policies are key objects in our analysis. We define them formally below, in a slightly generalized form. The history $\mathtt{hist}_{t}$ can be interpreted as a signal that is jointly distributed with state $\psi^{*}$. Indeed, fixing the mechanism in the previous rounds and assuming that the agents in these rounds report truthfully, the prior $\mathcal{P}$ on $\psi^{*}$ determines a joint distribution $\mathcal{Q}^{(t)}$ for the pair $(\psi^{*},\mathtt{hist}_{t})$. For the subsequent definitions, we replace the history with an abstract signal jointly distributed with $\psi^{*}$.
@@
As a shorthand, the Bayesian-expected utility of policy $\sigma$ under truthful reporting is
@@
This guarantee is “folklore”, see Appendix E for a proof.
@@ 3.1 Benchmark
We express our results in terms of a benchmark which encapsulates the optimal dependence on the prior. To this end, we consider recommendation policies that input the true state ($\mathtt{sig}=\psi^{*}$); we call them state-aware. We optimize among all BIR, state-aware policies $\sigma$ so as to minimize the error upper bound (2.6). More precisely, we’d like to minimize $\min_{a}1/\sigma_{a}(\psi^{*})$ in the worst case over adversaries $\mathtt{adv}$, i.e., in the worst case over all possible realizations of $\psi^{*}$.
@@ 3.2 Our mechanism: Main stage
To recap, our mechanism uses $(\epsilon,\overline{\psi})\text{-$\mathtt{greedy}$}$ recommendation policy in all rounds, for some $\epsilon>0$. Our guarantee is conditional on collecting enough samples in the warm-up stage.
@@
To complete the BIR proof, $U_{\mathcal{P}}(\sigma^{*})\geq U_{\mathtt{out}}$ by Lemma 3.2, since $2\epsilon=n/\mathtt{bench}(\mathcal{P})$.
@@ 3.3 Our mechanism: Warm-up stage
We rely on prior work on incentivized exploration to guarantee warm-up samples. Restated in our notation, the relevant guarantee is as follows:
@@
Thus, sufficient warm-up stage for Theorem 3.5 lasts $T_{0}=M_{\mathcal{P}}\cdot N_{\mathcal{P}}$ rounds, and consists of $N_{\mathcal{P}}$ consecutive runs of the mechanism from Lemma 3.6(a). The expression for $M_{\mathcal{P}}$ is somewhat tedious, and not essential for this paper. [^21] However, let us present it for $n=2$ arms, for the sake of completeness. Consider $\Delta_{\psi}:=U_{\psi}(1)-U_{\psi}(2)$, the “gap” between the arms’ expected utilities. W.l.o.g., assume that $\operatornamewithlimits{\mathbb{E}}_{\psi}\left[\,\Delta_{\psi}\,\right]\geq 0$, i.e., arm $1$ is (weakly) preferred initially. Let $X_{(k)}:=\operatornamewithlimits{\mathbb{E}}_{\psi}\left[\,-\Delta_{\psi}\mid S_{(k)}\,\right]$, where $S_{(k)}$ is a tuple of $k$ independent reward samples from arm $1$. Then, suppose there exists some finite $k=k_{\mathcal{P}}$ such that $\Pr\left[\,X_{(k)}>0\,\right]>0$ (if not, set $M_{\mathcal{P}}=\infty$). Let $Y=X_{(k_{\mathcal{P}})}$. Then $M_{\mathcal{P}}=(n+1)\max\left(\,k_{\mathcal{P}},\;\frac{\operatornamewithlimits{\mathbb{E}}_{\psi}\left[\,\Delta_{\psi}\,\right]}{\operatornamewithlimits{\mathbb{E}}\left[\,Y\mid Y>0\,\right]\;\Pr[Y>0]}\,\right).$
@@ 3.4 Lower bound
Thus, Theorem 3.7 allows per-round policies that change over time and can themselves be jointly chosen at random. In rounds $t\not\in S$ the policies can depend on all past observations. Importantly, this includes “multi-stage” mechanisms that partition the time into a constant number of contiguous “stages” and are non-data-adaptive within each (possibly with a data-adaptive warm-up stage of duration, say, $<T/2$). Most medical trials in practice fit this model, see Footnote 10.
@@ 4.1 Overview
We obtain matching upper and lower bounds like in Section 3, but with a new benchmark, mechanism, and analysis. This novelty is necessitated by the BIC condition which binds the types together. [^22] Indeed, without private types one can use mechanism from Section 3 tailored to each particular type. We define a new benchmark, denoted $\mathtt{bench}(\mathcal{P},\Theta)$, which generalizes (3.1) to a given set $\Theta$ of agent types. Compared to the expression in (3.1), this benchmark requires all BIC and BIR per-round policies under the $\inf$, and takes the $\max$ over $\Theta$ (as well as over all states). Our mechanism attains the same error bound as in Theorem 3.5 relative to $\mathtt{bench}(\mathcal{P},\Theta)$, with a similar but much more technical sufficient condition for BIR. The lower bound carries over from Theorem 3.7, replacing $\mathtt{bench}(\mathcal{P})$ with $\mathtt{bench}(\mathcal{P},\Theta)$.
@@ 4.2 Model: heterogeneous agents
Given a BIC mechanism, it is a (weakly) best response for each agent to report the private type truthfully, assuming the previous agents did so. We assume then that all agents report truthfully.
@@ 5 Heterogeneous agents with estimated type frequencies
We provide upper and lower bounds relative to this benchmark. Our mechanism attains a strong guarantee when the estimates $\widetilde{F}_{\mathtt{type}}$ are not too far from $F_{\mathtt{type}}$. When $\widetilde{F}_{\mathtt{type}}=F_{\mathtt{type}}$, this guarantee are optimal, much like in Theorem 3.7, in the worst case over all adversaries with a given $F_{\mathtt{type}}$.
@@ Appendix A Simple numerical examples
Our participation-incentivizing RCT design from Section 3 works out as follows. Leveraging Lemma 3.2, consider $\epsilon^{*}$, the largest $\epsilon$ such that $(\epsilon,\psi^{*})\text{-$\mathtt{greedy}$}$ policy is BIR. We claim $\epsilon^{*}=\nicefrac{{8}}{{9}}$. Indeed, note that $\epsilon^{*}$ is is the unique $\epsilon$ such that the Bayesian-expected utility of $(\epsilon,\psi^{*})\text{-$\mathtt{greedy}$}$, call it $U(\epsilon)$, equals the outside option $U_{\mathtt{out}}$. And one can check that $U(\nicefrac{{8}}{{9}})=\nicefrac{{2}}{{3}}=U_{\mathtt{out}}$. (One way to do that is to sum over the four possible states, as per Footnote 28, and for each state calculate the expected utility of the policy; we omit the easy details.)
@@
Now let us calculate $N_{\mathcal{P}}$, the number of samples required for the warm-up stage. The prior-expected utility of the best arm is $U_{\mathcal{P}}(a^{*}())=\nicefrac{{4}}{{3}}$ (the calculation is similar to that for $U(\epsilon^{*})$.) So, the “prior gap” (see Eq. (3.2)) is $\mathtt{gap}(\mathcal{P})=\nicefrac{{4}}{{3}}-\nicefrac{{5}}{{12}}=\nicefrac{{11}}{{12}}$. Therefore, according to Eq. 3.3, we have $\alpha=\nicefrac{{22}}{{27}}$ and $N_{P}<144$. In other words, a warm-up stage with $144$ independent samples of each arm suffices to guarantee incentive-compatibility for the main stage. Recall that these samples may also be obtained exogenously, e.g., via paid volunteers.

### v11

@@ 3 Our model: incentivized RCT (homogeneous agents)
We focus on homogeneous agents, i.e., agents with the same “type”. We consider a randomized controlled trial (RCT) with some patients/users, respectively interpreted as a mechanism and agents. The mechanism interacts with the agents as a multi-armed bandit algorithm, but optimizes a very different statistical objective, and operates under economic constraints to incentivize participation. Accordingly, we split the model description into three subsections: the bandit model for the interaction protocol the statistical model for its performance objective, and the economic model for the agents’ incentives.
@@ 3.2 Statistical model: the performance objective
The mechanism should estimate the average score of each arm without any assumptions on the adversary, i.e., on how the outcomes are generated. Not relying on modeling assumptions is a standard desideratum in randomized controlled trials.
@@
We focus on the following two scenarios. First, $f_{\mathtt{adv}}(a)$ is the frequency of a given outcome $\omega^{*}\in\Omega$ in the main stage; then the scoring function is $f(\omega)={\bf 1}_{\left\{\,\omega=\omega^{*}\,\right\}}$. Second, if $f(\omega)$ is the outcome’s value (for the agent and/or the society), then $f_{\mathtt{adv}}(a)$ is the average value of arm $a$. In particular, the average treatment effect can be expressed as $f_{\mathtt{adv}}(\text{treatment})-f_{\mathtt{adv}}(\text{control})$. We keep $f$ abstract throughout to handle both scenarios in a uniform way.
@@ 3.3 Economic model: agents, beliefs and incentives
The mechanism must incentivize participation, i.e., satisfy Individual Rationality (IR). We require Bayesian IR in expectation over the agents’ beliefs over the adversaries. [^16] We invoke agents’ beliefs because satisfying the IR property for every adversary (a.k.a. ex-post IR) is clearly impossible, except for a trivial mechanism that always chooses the agent’s outside option. Crucially, we posit that agents’ beliefs are over stochastic adversaries, i.e., the agents believe that a given treatment yields the same outcome distribution in all agents. 171717 We called such beliefs time-invariant in the Introduction. These beliefs are misspecified, meaning that their support might not include the actual adversary (which need not be stochastic). See Related Work for background. All agents share the same belief, a common simplification in economic theory. Such beliefs represent the current medical knowledge relevant to the clinical trial, or the learnings from user interactions collected by the online platform (which the platform may choose to share with its users).
@@ 3.4 Preliminaries
Per-round recommendation policy. A single round $t$ of our mechanism can be interpreted as a stand-alone “recommendation policy” which inputs the $\mathtt{hist}_{t}$, and outputs a recommended arm. Such policies are key objects in our analysis. We define them formally below.
We focus on a hypothetical world that operates according to agents’ beliefs, whereby we have stochastic adversaries and the true state $\psi^{*}$ is drawn from the prior $\mathcal{P}$. The history $\mathtt{hist}_{t}$ can be interpreted as a signal that is jointly distributed with state $\psi^{*}$. Indeed, the prior and the mechanism in the previous rounds determine a joint distribution $\mathcal{Q}^{(t)}$ for the pair $(\psi^{*},\mathtt{hist}_{t})$.
We will need a more general definition in which we replace the history with an abstract signal jointly distributed with $\psi^{*}$. (The joint signal-state distribution is determined by the prior and the signal-generating process; we leave the latter outside of our definition.)
@@
As a shorthand, the Bayesian-expected utility of policy $\sigma$ is
@@
This guarantee is “folklore”, see Appendix E for a proof. This is the statistical tool we use to analyze our mechanisms. For the lower bounds, we use a somewhat non-standard tool which applies to arbitrary counterfactual estimators (Eq. (4.5) and Appendix D).
@@ 4.1 Benchmark
We express our results in terms of a benchmark which focuses on a single round and encapsulates the optimal dependence on the prior. To this end, we consider recommendation policies that input the true state ($\mathtt{sig}=\psi^{*}$); we call them state-aware. We optimize among all BIR, state-aware policies $\sigma$ so as to minimize the error upper bound (3.6). More precisely, we’d like to minimize $\max_{a}1/\sigma_{a}(\psi^{*})$ in the worst case over adversaries $\mathtt{adv}$, i.e., in the worst case over all possible realizations of $\psi^{*}$.
@@
Note that the right-hand side of (4.2) also depends on the prior $\mathcal{P}$ through the definition of BIR.
@@ 4.2 Our mechanism: Main stage
To recap, our mechanism uses $(\epsilon,\overline{\psi})\text{-$\mathtt{greedy}$}$ recommendation policy in all rounds, for some $\epsilon>0$. (Unlike the standard epsilon-greedy algorithm, the estimated best arm is fixed throughout, by the choice of $\overline{\psi}$, rather than adjusted as more data comes in.) Our guarantee is conditional on collecting enough samples in the warm-up stage.
@@
To complete the BIR proof, since $2\epsilon=n/\mathtt{bench}(\mathcal{P})$, policy $\sigma^{*}$ is precisely a $(n/\mathtt{bench}(\mathcal{P}),\psi^{*})$-greedy policy. According to Lemma 4.2, this policy is BIR and $U_{\mathcal{P}}(\sigma^{*})\geq U_{\mathtt{out}}$.
@@ 4.3 Our mechanism: Warm-up stage
We leverage prior work on incentivized exploration to guarantee warm-up samples. Restated in our notation, the relevant guarantee is as follows:
@@
While Lemma 4.6(a) is trivial as stated above, Mansour et al., (2020) provide an explicit mechanism for part (a) and an explicit expression for $M_{\mathcal{P}}$. The latter is somewhat tedious, and not essential for this paper. [^20] However, let us present it for $n=2$ arms, for the sake of completeness. Consider $\Delta_{\psi}:=U_{\psi}(1)-U_{\psi}(2)$, the “gap” between the arms’ expected utilities. W.l.o.g., assume that $\operatornamewithlimits{\mathbb{E}}_{\psi}\left[\,\Delta_{\psi}\,\right]\geq 0$, i.e., arm $1$ is (weakly) preferred initially. Let $X_{(k)}:=\operatornamewithlimits{\mathbb{E}}_{\psi}\left[\,-\Delta_{\psi}\mid S_{(k)}\,\right]$, where $S_{(k)}$ is a tuple of $k$ independent reward samples from arm $1$. Then, suppose there exists some finite $k=k_{\mathcal{P}}$ such that $\Pr\left[\,X_{(k)}>0\,\right]>0$ (if not, set $M_{\mathcal{P}}=\infty$). Let $Y=X_{(k_{\mathcal{P}})}$. Then $M_{\mathcal{P}}=(n+1)\max\left(\,k_{\mathcal{P}},\;\frac{\operatornamewithlimits{\mathbb{E}}_{\psi}\left[\,\Delta_{\psi}\,\right]}{\operatornamewithlimits{\mathbb{E}}\left[\,Y\mid Y>0\,\right]\;\Pr[Y>0]}\,\right).$ Thus, a sufficient warm-up stage for Theorem 4.5 lasts $T_{0}=M_{\mathcal{P}}\cdot N_{\mathcal{P}}$ rounds, and consists of $N_{\mathcal{P}}$ consecutive runs of the mechanism from Lemma 4.6(a).
Remark 4.7.
While Mansour et al., (2020) assumes stochastic adversaries, this suffices because the BIR property is required only for time-invariant beliefs. So, the mechanism from Lemma 4.6(a) is BIR regardless of whether the actual adversary is stochastic. The mechanism samples all arms in $M_{\mathcal{P}}$ rounds deterministically, regardless of the observed outcomes.
@@ 4.4 Lower bound
Thus, Theorem 4.8 allows per-round policies that change over time and can themselves be jointly chosen at random. In rounds $t\not\in S$, the policies can depend on all past observations. Importantly, this includes “multi-stage” mechanisms that partition the time into a constant number of contiguous “stages” and are non-data-adaptive within each (possibly with a data-adaptive warm-up stage of duration, say, $<T/2$). Most medical trials in practice fit this model, see Footnote 9.
The proof plan is as follows. First, we derive a statistical lower bound for an arbitrary estimator (Eq. (4.5)). Second, we focus on a particular “bad state” and “bad arm” for the benchmark. Third, we construct a “bad” adversary, using this bad state/arm pair. The final computation demonstrates that this adversary gives a large enough error. We flesh out this plan below.
@@ 5.1 Overview
We obtain matching upper and lower bounds like in Section 4, but with a new benchmark, mechanism, and analysis. This novelty is necessitated by the Bayesian Incentive Compatibility (BIC) constraint which incentivizes truthful reporting of private types and binds the types together. [^21] Indeed, without private types one can use mechanism from Section 4 tailored to each particular type. We define a new benchmark, denoted $\mathtt{bench}(\mathcal{P},\Theta)$, which generalizes (4.1) to a given set $\Theta$ of agent types. Compared to the expression in (4.1), this benchmark requires all BIC and BIR per-round policies under the $\inf$, and takes the $\max$ over $\Theta$ (as well as over all states). Our mechanism attains the same error bound as in Theorem 4.5 relative to $\mathtt{bench}(\mathcal{P},\Theta)$, with a similar but much more technical sufficient condition for BIR. The lower bound carries over from Theorem 4.8, replacing $\mathtt{bench}(\mathcal{P})$ with $\mathtt{bench}(\mathcal{P},\Theta)$.
@@ 5.2 Model: heterogeneous agents
Given a BIC mechanism, it is a (weakly) best response for each agent to report the private type truthfully, assuming the previous agents did so. We assume that all agents report truthfully.
Remark 5.6.
Private type elicitation does not have to be BIC. However, mechanisms in our model need to interpret (mis)reported private types, one way or another, and the BIC constraint is one standard way to proceed. Note that restricting to BIC mechanisms and truthful reporting is without loss of generality according to Myerson’s revelation principle (Myerson,, 1981).
@@ 5.3 Our results
Remark 5.11.
The structure of $\sigma^{\mathtt{opt}}(\psi^{*},\cdot)$, the optimal state-aware policy, is not essential for our analysis. Nonetheless, we state it below for the sake of clarity. Consider a “menu” of per-type optimal policies, which are the $(\epsilon,\psi^{*})$-greedy policies with tight BIR constraints for the respective type. Then each type is mapped to the policy from this menu that maximizes its utility. This mapping ensures truthful reporting of private types.
@@ 6 Heterogeneous agents with estimated type frequencies
This benchmark is no larger than the worst-case benchmark $\mathtt{bench}(\mathcal{P},\Theta)$, and can be much smaller; we provide a simple numerical example in Appendix A.
We provide upper and lower bounds relative to $\mathtt{bench}(\mathcal{P},F_{\mathtt{type}})$. Our mechanism attains a strong guarantee when the estimated type frequencies $\widetilde{F}_{\mathtt{type}}$ are not too far from $F_{\mathtt{type}}$. When $\widetilde{F}_{\mathtt{type}}=F_{\mathtt{type}}$, this guarantee is optimal, much like in Theorem 4.8, in the worst case over all adversaries with a given $F_{\mathtt{type}}$.
@@ Appendix A Simple numerical examples
Our participation-incentivizing RCT design from Section 4 works out as follows. Leveraging Lemma 4.2, consider $\epsilon^{*}$, the largest $\epsilon$ such that $(\epsilon,\psi^{*})\text{-$\mathtt{greedy}$}$ policy is BIR. We claim $\epsilon^{*}=\nicefrac{{8}}{{9}}$. Indeed, note that $\epsilon^{*}$ is is the unique $\epsilon$ such that the Bayesian-expected utility of $(\epsilon,\psi^{*})\text{-$\mathtt{greedy}$}$, call it $U(\epsilon)$, equals the outside option $U_{\mathtt{out}}$. And one can check that $U(\nicefrac{{8}}{{9}})=\nicefrac{{2}}{{3}}=U_{\mathtt{out}}$. (One way to do that is to sum over the four possible states, as per Footnote 27, and for each state calculate the expected utility of the policy; we omit the easy details.)
@@
Now let us calculate $N_{\mathcal{P}}$, the sufficient number of samples for the warm-up stage, as per Eq. (4.2) in our analysis. [^28] While the expression in Eq. (4.2) has an appealing “interpretability”, characterizing $N_{\mathcal{P}}$ in terms of the prior gap $\mathtt{gap}(\mathcal{P})$ and the benchmark $\mathtt{bench}(\mathcal{P})$, we suspect one can improve over Eq. (4.2) with a more elaborate analysis. The prior-expected utility of the best arm is $U_{\mathcal{P}}(a^{*}())=\nicefrac{{4}}{{3}}$ (the calculation is similar to that for $U(\epsilon^{*})$.) So, the “prior gap” (see Eq. (4.2)) is $\mathtt{gap}(\mathcal{P})=\nicefrac{{4}}{{3}}-\nicefrac{{5}}{{12}}=\nicefrac{{11}}{{12}}$. Therefore, according to Eq. 4.3, we have $\alpha=\nicefrac{{22}}{{27}}$ and $N_{P}<144$. In other words, a warm-up stage with $144$ independent samples of each arm suffices to guarantee incentive-compatibility for the main stage. Recall that these samples may also be obtained exogenously, e.g., via paid volunteers.